OpenStudy (aravindg):

if the earth rotates x times faster than its present speed about its axis in rder that the bodies lying on its equator just fly off into space ,then x nearly equals??

OpenStudy (jamesj):

So what is the centripetal force? mv^2/r. You need that to be greater than the gravitational force, mg.

OpenStudy (jamesj):

Alternatively, write centripetal force as \[ F_c = m\omega^2 r \] Now use that to find the angular velocity \( \omega \) required for things to spin off the earth and compare that with the earth's current angular velocity.

OpenStudy (aravindg):

well i tried like this g'=g(1-rw^2/g cos^2 theta)

OpenStudy (jamesj):

at the equator theta = 0.

OpenStudy (aravindg):

ys

OpenStudy (aravindg):

g'=g(1-rw^2/g)

OpenStudy (jamesj):

The equation you've written down is the acceleration of a body experiencing both forces, as the centripetal force acts against gravity here \[ F_g - F_c = mg - m\omega^2 r = m(g - \omega^2r) = m.g.(1 - \omega^2r/g) \]

OpenStudy (jamesj):

So whatever you like. You want to find the \( \omega \) such that that acceleration, what you call g', is equal to zero.

OpenStudy (aravindg):

wel according to my equation rw^2/g should become equal to 1 is it?

OpenStudy (jamesj):

Yes. The force is zero and g' = 0 precisely when \[ frac{r \omega^2}{g} = 1 \] Equivalently, when \[ r\omega^2 = g \] or \[ \omega = \sqrt{g/r} \]

OpenStudy (jamesj):

* \[ \frac{r\omega^2}{g} = 1 \]

OpenStudy (aravindg):

and james u made a typo its centrifugal force not centripetal

OpenStudy (jamesj):

yes, fine.

OpenStudy (aravindg):

so where does this x come?

OpenStudy (aravindg):

i mean we have to compare with present omega and xomega

OpenStudy (jamesj):

Find the \( \omega \) for which things would fly off at the equator. And compare it with the current angular velocity of the earth, call that \( \omega_0 \). Then by definition of x, "how many times faster than the earth spins now", \[ x = \frac{\omega}{\omega_0} \]

OpenStudy (aravindg):

for that we need value of w.do u knw tht?

OpenStudy (jamesj):

C'mon. You can figure out omega from the formula you wrote down above. And omega_0, the current angular velocity of the earth you can figure out!! How long does it take the earth to complete one rotation, \( 2\pi \) ?

OpenStudy (aravindg):

hmm.7.27*10^-5

OpenStudy (aravindg):

:)

OpenStudy (jamesj):

right.

OpenStudy (aravindg):

wat abt radius isnt it 6400000m

OpenStudy (aravindg):

wow!!i got the final answer as 17!

OpenStudy (jamesj):

Yes, I got that too.

OpenStudy (aravindg):

yey thanks for help

OpenStudy (aravindg):

one last qn and i am ovr with gravitation

OpenStudy (aravindg):

ready?